OpenStudy (anonymous):

differentiate y=ln2x/sqrtx using quotient rule

OpenStudy (anonymous):

the quotient rule says that  here, f(x) = ln(2x) and g(x) = sqrt(x) f'(x) = 1/x and g'(x) = 1/(2sqrt(x)) now you should just be able to plug it in